# The dimension and Bose distance of certain primitive BCH codes

### v2

@@ V The Dimension of $\mathcal{C}_{(q,m,\delta,b)}$ for $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1$
$(\delta^{{}^{\prime}}_{j_{\delta}+r_{\delta}-1},\ldots,\delta^{{}^{\prime}}_{0})\leq(\delta_{h+j_{\delta}},\ldots,\delta_{h-r_{\delta}+1})=(\delta^{{}^{\prime}}_{h+j_{\delta}},\ldots,\delta^{{}^{\prime}}_{h-r_{\delta}+1}).$ (69)
@@
Suppose that $m\geq 4$ is an even integer, and $\delta\in[q^{h},q^{\lfloor({2m-1})/{3}\rfloor+1})$ is an integer. Let $j_{\delta}$ be the integer such that $q^{h+j_{\delta}}\leq\delta<q^{h+j_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ denote the $q$-adic expansion of $\delta$, and let $r_{\delta}$ be the smallest integer in $[-j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$. Define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})},\\ \hat{\delta}+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\\ &\hbox{ and }\delta_{h-r_{\delta}}\neq q-1,\\ \hat{\delta}+2,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\\ &\hbox{ and }\delta_{h-r_{\delta}}=q-1.\\ \end{cases}$ (70)
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})},\\ \hat{\delta}+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\\ &\hbox{ and }\delta_{h-r_{\delta}}\neq q-1,\\ \hat{\delta}+2,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\\ &\hbox{ and }\delta_{h-r_{\delta}}=q-1.\\ \end{cases}$ (71)
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}\\ &\hbox{ and }q\nmid\delta,\\ \delta+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}\\ &\hbox{ and }q\mid\delta,\\ \hat{\delta},&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}.\\ \end{cases}$ (72)
@@
$(\mathbb{0}_{h-j_{\delta}-1},\delta_{j_{\delta}},\ldots,\delta_{0},\mathbb{0}_{h-j_{\delta}-1},\delta_{j_{\delta}},\ldots,\delta_{0})$
@@
Subcase 2. If $\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell-h}<\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}.$ By an argument similar to that in Case 1 of the proof of Theorem 6, we first conclude that any integer $\delta^{{}^{\prime}}\in\left[\delta,\hat{\delta}\right]$ is not a coset leader. Note that $\delta_{h}>0$ and $\hat{\delta}=\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}$ when $r_{\delta}=0$. Applying Corollary 2, one can easily verify that $\hat{\delta}\in\mathcal{H}$. This implies that $\hat{\delta}$ is a coset leader. Consequently, we have $d_{B}(\mathcal{C}_{(q,m,\delta)})=\hat{\delta}$. ∎
An $[n,k,d]$ code over $\mathbb{F}_{q}$ is called optimal if there is no $[n,k^{\prime},d]$ linear code with $k^{\prime}\geq k$, or $[n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. We provide examples of narrow-sense BCH codes discussed in this paper in Tables III and III. The dimensions $k$ and Bose distances $d_{B}$ of the BCH codes presented in these tables are calculated using the formulas from Theorems 3, 5, 6, and 7, with all parameters verified using Magma. We compare these codes with the Database and find that they are either optimal or match the parameters of the best-known linear codes. We also utilized Magma to calculate the minimum distance of the codes. For codes whose minimum distance could not be determined by Magma within a reasonable time, we denote this with ”–”. Notably, the Bose distance and minimum distance of these codes are equal.
Table II: Example of binary BCH code $\mathcal{C}_{(q,m,\delta)}$
for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$
$q$ $m$ $\delta$ $n$ $k$ $d_{B}$ $d$ Optimality
@@
Table III: Example of nonbinary BCH code $\mathcal{C}_{(q,m,\delta)}$
for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$
$q$ $m$ $\delta$ $n$ $k$ $d_{B}$ $d$ Optimality
@@
VII Dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ with $\delta=aq^{h+k}+b$.
As a special case of our main results in the above sections, we give the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta=q^{h+k}+b$ with $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, $a\in[1,q-1]$ and $b\in\left[1,q^{m-h-k}\right]$ in the following corollaries.
@@
Let $m\geq 5$ be an odd intger and $k\in[1,\lfloor(2m-1)/3\rfloor-h]$ be an integer. If $\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[1,q^{h-k+1}]$, then
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}n-mN(\delta)+ma^{2}(q-1)q^{2k-3}\left[(q-1)k+1\right],&\hbox{if }b>aq^{2k-1},\\ n-maq^{h+k-1}(q-1)+ma(q-1)q^{2k-3}\left[a(q-1)k+a-q\right],&\hbox{if }b\leq aq^{2k-1}.\end{cases}$ (74)
@@
$(\mathbb{0}_{h-k},a,\mathbb{0}_{2k-1},\delta_{h-k},\ldots,\delta_{0})$
@@
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-3},&\hbox{if }i\in[-k+2,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-2},&\hbox{if }i=k\hbox{ or }-k+1.\end{cases}$
@@
$\begin{split}\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)&=\sum\limits_{i=-k+1}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)\\ &=a(a-1)(q-1)q^{2k-2}+(a^{2}-1)(k-1)(q-1)^{2}q^{2k-3}.\end{split}$
@@
$f(\delta)=\begin{cases}a^{2}(q-1)q^{2k-3}\left[k(q-1)+1\right],&\hbox{if }b>aq^{2k-1},\\ a(q-1)q^{2k-3}\left[ak(q-1)+a-q\right]+N(b),&\hbox{if }b\leq aq^{2k-1}.\end{cases}$
@@
$d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}aq^{h}+a,&\hbox{if }k=0\hbox{ and }b\leq a,\\ \delta,&\hbox{if }k=0\hbox{ and }b>a,\\ \delta,&\hbox{if }k\geq 1\hbox{ and }b>aq^{2k},\\ aq^{h+k}+aq^{2k}+1,&\hbox{if }k\geq 1\hbox{ and }b\leq aq^{2k}.\end{cases}$ (75)
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}n-maq^{h-1}(q-1)+\frac{1}{2}m(a-1)^{2},&\hbox{if }k=0\hbox{ and }b\leq a,\\ n-mN(\delta)+\frac{1}{2}ma^{2},&\hbox{if }k=0\hbox{ and }b>a,\\ n-mN(\delta)+ma^{2}(q-1)q^{2k-2}\left[\left(k-\frac{1}{2})(q-1\right)+q\right],&\hbox{if }k\geq 1\hbox{ and }b>aq^{2k},\\ n-maq^{h+k-1}(q-1)+maq^{2k-2}(q-1)\left[a(k-\frac{1}{2})(q-1)+(a-1)q\right],&\hbox{if }k\geq 1\hbox{ and }b\leq aq^{2k}.\end{cases}$ (76)
@@
$(\mathbb{0}_{h-k-1},a,\mathbb{0}_{2k},\delta_{h-k-1},\ldots,\delta_{0})$
@@
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i}+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-2},&\hbox{if }i\in[-k+1,0)\cup(0,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-1},&\hbox{if }i=k\hbox{ or }-k,\\ \frac{1}{2}(a-1)q^{k-1}(q-1)+\frac{1}{2}(a^{2}-1)q^{2k-2}(q-1)^{2},&\hbox{if }i=0.\end{cases}$
@@
$\begin{split}\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i}+1)&=\sum\limits_{i=-k}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i}+1)\\ &=(a^{2}-1)(q-1)^{2}q^{2k-2}(k-\frac{1}{2})+a(a-1)(q-1)q^{2k-1}+\frac{1}{2}(a-1)q^{k-1}(q-1).\end{split}$
@@
$\widetilde{f}(\delta)=\begin{cases}a^{2}(q-1)q^{2k-2}\left[\left(k-\frac{1}{2}\right)(q-1)+q\right]+\frac{1}{2}aq^{k-1}(q-1),&\hbox{if }k\geq 1\hbox{ and }b>aq^{2k},\\ aq^{2k-2}(q-1)\left[a(k-\frac{1}{2})(q-1)+(a-1)q\right]+\frac{1}{2}aq^{k-1}(q-1)+N(b),&\hbox{if }k\geq 1\hbox{ and }b\leq aq^{2k},\end{cases}$

### v3

@@ II Preliminaries
$\begin{split}(\delta^{{}^{\prime}}_{j_{\delta}+r_{\delta}-1},\ldots,\delta^{{}^{\prime}}_{0})&\leq(\delta_{h+j_{\delta}},\ldots,\delta_{h-r_{\delta}+1})\\ &=(\delta^{{}^{\prime}}_{h+j_{\delta}},\ldots,\delta^{{}^{\prime}}_{h-r_{\delta}+1}).\end{split}$ (66)
@@
Suppose that $m\geq 4$ is an even integer, and $\delta\in[q^{h},q^{\lfloor({2m-1})/{3}\rfloor+1})$ is an integer. Let $j_{\delta}$ be the integer such that $q^{h+j_{\delta}}\leq\delta<q^{h+j_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ denote the $q$-adic expansion of $\delta$, and let $r_{\delta}$ be the smallest integer in $j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$. Define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
@@
[…]
@@
[…]
@@
[…]
@@
$(\mathbf{0}_{h-j_{\delta}-1},\delta_{j_{\delta}},\ldots,\delta_{0},\mathbf{0}_{h-j_{\delta}-1},\delta_{j_{\delta}},\ldots,\delta_{0})$
@@
[…]
An $[n,k,d]$ code over $\mathbb{F}_{q}$ is called optimal if there is no $n,k^{\prime},d]$ linear code with $k^{\prime}>k$, or $n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. We provide examples of narrow-sense BCH codes discussed in this paper in Tables II and III. The dimensions $k$ and Bose distances $d_{B}$ of the BCH codes presented in these tables are calculated using the formulas from Theorems 3, 5, 6, and 7, with all parameters verified using Magma. We compare these codes with the Database and find that they are either optimal or match the parameters of the best-known linear codes. We also utilized Magma to calculate the minimum distance of the codes. For codes whose minimum distance could not be determined by Magma within a reasonable time, we denote this with ”–”. Notably, the Bose distance and minimum distance of these codes are equal.
Table II: Example of binary BCH code $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$
$q$ $m$ $\delta$ $n$ $k$ $_{B} dd$ Optimality
@@
Table III: Example of nonbinary BCH code $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/3\rfloor+1}]$
$q$ $m$ $\delta$ $n$ $k$ $_{B} dd$ Optimality
@@
[…]
As a special case of our main results in the above sections, we give the dimension and Bose distance of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $\delta=aq^{h+k}+b$ with $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, $a\in[1,q-1]$ and $b\in\left[1,q^{m-h-k}\right]$ in the following corollaries.
@@
Let $m\geq 5$ be an odd integer and $k\in[1,\lfloor(2m-1)/3\rfloor-h]$ be an integer. If $\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[1,q^{h-k+1}]$, then
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\\ \begin{cases}\left(\begin{aligned} &n-mN(\delta)\\ &+ma^{2}(q-1)q^{2k-3}\left[(q-1)k+1\right]\end{aligned}\right),&\hskip-7.0pt\hbox{if }b>aq^{2k-1},\\ \left(\begin{aligned} &n-maq^{h+k-1}(q-1)\\ &-ma(q-1)q^{2k-2}\\ &+ma^{2}(q-1)q^{2k-3}\left[(q-1)k+1\right]\end{aligned}\right),&\hskip-7.0pt\hbox{if }b\leq aq^{2k-1}.\end{cases}$ (71)
@@
$(\mathbf{0}_{h-k},a,\mathbf{0}_{2k-1},\delta_{h-k},\ldots,\delta_{0})$
@@
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-3},\\ \quad\quad\hbox{if }i\in[-k+2,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-2},\\ \quad\quad\hbox{if }i=k\hbox{ or }-k+1.\end{cases}$
@@
[…]
$=$ $\sum\limits_{i=-k+1}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)$
$=$ $a(a-1)(q-1)q^{2k-2}+(a^{2}-1)(k-1)(q-1)^{2}q^{2k-3}.$
@@
$f(\delta)=\begin{cases}a^{2}(q-1)q^{2k-3}\left[k(q-1)+1\right],\hbox{ if }b>aq^{2k-1},\\ a(q-1)q^{2k-3}\left[ak(q-1)+a-q\right]+N(b),\\ \quad\quad\hbox{if }b\leq aq^{2k-1}.\end{cases}$
@@
$d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}aq^{h}+a,&\hbox{if }k=0,b\leq a,\\ \delta,&\hbox{if }k=0,b>a,\\ \delta,&\hbox{if }k\geq 1,b>aq^{2k},\\ aq^{h+k}+aq^{2k}+1,&\hbox{if }k\geq 1,b\leq aq^{2k}.\end{cases}$ (72)
@@
$\begin{split}&\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\\ &\begin{cases}n-maq^{h-1}(q-1)+\frac{1}{2}m(a-1)^{2},&\hbox{if }k=0,b\leq a,\\ n-mN(\delta)+\frac{1}{2}ma^{2},&\hbox{if }k=0,b>a,\\ \left(\begin{aligned} &ma^{2}(q-1)^{2}q^{2k-2}\left(k-\textstyle\frac{1}{2}\right)\\ &+ma^{2}(q-1)q^{2k-1}\\ &+n-mN(\delta)\end{aligned}\right),&\hbox{if }k\geq 1,b>aq^{2k}\\ \left(\begin{aligned} &ma^{2}(q-1)^{2}q^{2k-2}\left(k-\textstyle\frac{1}{2}\right)\\ &+ma(a-1)(q-1)q^{2k-1}\\ &+n-maq^{h+k-1}(q-1)\end{aligned}\right),&\hbox{if }k\geq 1,b\leq aq^{2k}.\end{cases}\end{split}$ (73)
@@
$(\mathbf{0}_{h-k-1},a,\mathbf{0}_{2k},\delta_{h-k-1},\ldots,\delta_{0})$
@@
[…]
@@
$\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i}+1)$ $=$ $\sum\limits_{i=-k}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i}+1)$
$=$ $(a^{2}-1)(q-1)^{2}(k-\frac{1}{2})q^{2k-2}$
$+\>a(a-1)(q-1)q^{2k-1}$
$+\>\frac{1}{2}(a-1)q^{k-1}(q-1).$
@@
$\widetilde{f}(\delta)=\begin{cases}\left(\begin{aligned} &a^{2}(q-1)^{2}q^{2k-2}\left(k-\textstyle\frac{1}{2}\right)\\ &+a^{2}(q-1)q^{2k-1}\\ &+\textstyle\frac{1}{2}aq^{k-1}(q-1)\end{aligned}\right),&\hskip-8.0pt\hbox{if }k\geq 1,b>aq^{2k},\\ \left(\begin{aligned} &a^{2}(q-1)^{2}q^{2k-2}(k-\textstyle\frac{1}{2})\\ &+a(a-1)(q-1)q^{2k-1}\\ &+\textstyle\frac{1}{2}aq^{k-1}(q-1)+N(b)\end{aligned}\right),&\hskip-8.0pt\hbox{if }k\geq 1,b\leq aq^{2k},\end{cases}$